# Thermal spectrometer for superconducting circuits

### v1

@@ Introduction
Figure 1: Principle of characterizing a resonator with a bolometer. a, Photons in the resonator (dark blue) are excited due to the input microwave tone $P_{\textrm{in}}$ that is injected to the nearby feedline (light blue). The photons escape via several loss channels, a fraction of them to a nearby capacitively coupled bolometer consisting of an absorber (red) and thermometer. By measuring temperature rise in the bolometer due to this photon leakage, the total loss rate ($\gamma_{\textrm{t}}=\gamma_{\textrm{f}}+\gamma_{\textrm{i}}+\gamma_{\textrm{b}}$) and resonance frequency of the resonator ($f_{0}$) can be determined. b, The system is the case of a driven resonator (a harmonic oscillator) coupled to several thermal baths that absorb the photons from the resonator. c, Equivalent lumped circuit. The bolometer is heated up due to the power $P_{\textrm{b}}$ transmitted to the absorber $R_{\textrm{b}}$. The magnitude of $P_{\textrm{b}}$ depends on the impedance of the other circuit elements $R,L,{\textrm{and}}~{}C$. $P_{\textrm{i}}$ is the power dissipated in the internal bath.
Detection in superconducting devices is a central key in the characterization and operation of the devices. For example, rf spectroscopy becomes a standard technique to measure qubit energy spectrum and its coherent interaction with microwave photons [16, 17, 18]. The study of a resonator and its loss mechanisms relies on the spectroscopy technique, either transmission ($S_{21}$) or reflection ($S_{11}$) measurement [19, 20]. Conventional rf spectroscopy, involving amplification of a microwave probe tone at different temperature stages followed by heterodyne detection at room temperature, requires costly electronics and cryogenic elements such as vector network analyzer (VNA), low-temperature amplifiers, circulators, and superconducting coaxial lines. These components are bulky and lossy. Precise cabling, grounding, and connection of the lines and components are crucial to avoid any microwave mismatch and leakage, that can result to parasitic resonances and suppression of the desired signals[21].
@@
A bolometer consists of an energy absorbing element and the associated thermometer detecting temperature changes due to energy input [22, 23]. Bolometric detection in a superconducting circuit offers advantages compared to conventional schemes of rf measurement. Using a bolometric detector, that can be placed at the millikelvin stage, the use of cryogenic elements can be avoided. A single-shot qubit readout has been demonstrated by using a bolometer that is located on a separate chip [24]. Without the use of low-temperature parametric amplifier, the authors could measure the qubit state by monitoring temperature in the bolometer. Another advantage is the broadband operation of the bolometer, while a typical rf setup is limited to narrow frequency range of $4-8~{}\textrm{GHz}$. A nano-bolometer has been used to calibrate the attenuation of a coaxial line in a broad frequency range [25]. A bolometric measurement of a high frequency Josephson ac current has been demonstrated [26] where the authors could detect ac current up to 100 GHz. Morever, the bolometer could reach small energy resolution as demonstrated by proximitized thermometer and graphene as the thermal detector [27, 28], and it is a potential candidate of a single photon detector.
@@
The principle of characterizing a superconducting resonator by an on-chip bolometer is shown in Fig. 1a. The resonator (dark blue) is coupled to a feedline (light blue) via capacitance $C_{\textrm{f}}$. The open-circuit end of the resonator (left) is coupled by capacitance $C_{\textrm{b}}$ to a bolometer that consists of an absorber (red) with resistance $R_{\textrm{b}}$ and the associated thermometer to detect temperature change due to input energy. A microwave tone $P_{\textrm{in}}$ is injected through the input feedline to excite photons in the resonator. The photons subsequently decay via several loss channels. Internal loss is mostly due to TLSs or QPs depending on power and temperature, and external losses are dominated by the proximity of the feedline and the bolometer itself. Thus, the total loss rate is $\gamma_{\textrm{t}}=\gamma_{\textrm{b}}+\gamma_{\textrm{i}}+\gamma_{\textrm{f}}$, where $\gamma_{\textrm{i}}$ is the internal loss rate, $\gamma_{\textrm{f}}$ is loss rate to the feedline, and $\gamma_{\textrm{b}}$ is loss rate to the bolometer. The system is the case of a driven resonator (a harmonic oscillator) coupled to several thermal baths, as shown in Fig.1b. The photons that are absorbed by the bolometer heat the resistor with power $P_{\textrm{b}}$, leading to temperature rise $T_{\textrm{b}}$. The power $P_{\textrm{b}}$ takes Lorentzian form as a function of the driving frequency around resonance $f_{0}$ with the linewidth $\gamma_{\textrm{t}}$. Due to the linearity of the resonator, $P_{\textrm{b}}$ is independent on bolometer temperature and takes a form
$P_{\mathrm{b}}(f)=\frac{1}{2}\frac{\gamma_{\textrm{f}}\gamma_{\textrm{b}}}{(f/f_{0}-1)^{2}+(\gamma_{\textrm{t}}/2)^{2}}P_{\mathrm{in}},$ (1)
a result that one can obtain both by circuit theory and open quantum system approach (see Supplementary Information I and II). By measuring the frequency dependence of $P_{\textrm{b}}$, which can be measured by a steady-state temperature measurement of the bolometer, the total-loss rate $\gamma_{\textrm{t}}$ and the resonance frequency $f_{0}$ of the resonator can thus be determined.
@@
Figure 3: Power measured by bolometer. a, Heating power $P_{\text{b}}$ and bolometer temperature $T_{\text{b}}$ due to photon leakage from the resonator. b, $P_{\text{b}}$ versus $f$ at different input powers varying from $-130~{}\mathrm{dBm}$ to $-95~{}\mathrm{dBm}$.
@@
Figure 5: Linewidth and quality factor. a, Total photon loss rate $\gamma_{\textrm{t}}$ versus $P_{\text{in}}$. b, Comparison between measured internal quality factor $Q_{\textrm{i}}$ (orange) and TLSs $Q_{\textrm{i,TLS}}$ (blue) and QPS $Q_{\textrm{i,QPS}}$ (red) model.
@@
The studied device consists of a quarter-wavelength ($\lambda/4$) coplanar waveguide (CPW) resonator capacitively probed by the feedline and the on-chip bolometer as shown in the device layout in Fig. 2(a). Both resonator and feedline are made of niobium (Nb) film on top of silicon (Si) substrate. We design $C_{\mathrm{f}}\sim C_{\mathrm{b}}\sim 20~{}\mathrm{fF}$, $R_{\mathrm{b}}$ is about $12~{}\Omega$, total inductance of the resonator is $L=83~{}\mathrm{pH}$, and total capacitance of the resonator is $C=333~{}\mathrm{fF}$. These values correspond to the quality factor of the feedline and bolometer as $Q_{\mathrm{f}}\sim 1681$ and $Q_{\mathrm{b}}\sim 1716$ respectively. The shorted end of the resonator (right) is shunted by inactive flux qubit consisting of a parallel aluminium (Al) line and Josephson junctions, tuned at zero magnetic flux thus effectively shunted only by the Al line [30, 31]. The open-circuit end of the resonator is capacitively connected to a copper (Cu) film functioning as the bolometer as shown in Fig. 2(b). In the experiment, we employ a pair of normal metal-insulator-superconductor (NIS) junctions in SINIS configuration to probe the temperature ($T_{\textrm{b}}$) in the Cu absorber (red), as shown in Fig. 2(c). The temperature rise manifests in the change of voltage $V_{\textrm{th}}$ accross the SINIS pair [23, 32]. Once the temperature is determined from the calibration conversion (see Supplementary Information III), the power can be calculated by electron phonon relation as
@@
We convert the measured $V_{\text{th}}$ to its corresponding temperature $T_{\text{b}}$ and power dissipation $P_{\text{b}}$. By using the conversion formula obtained from the bolometer calibration (see Supplementary Information III), we obtain $T_{\text{b}}$ around the resonance as shown in Fig  3a on the right axis. Although the measurement is carried out at 50 mK, $T_{\text{b}}$ saturates at around 130 mK. This saturation is dominated by heating from the environment with effective background power $P_{\text{e}}\sim 2~{}\text{fW}$ [32, 34]. The heating power $P_{\text{b}}$, converted from Eq.3, is plotted in Fig  3a on the left axis. Figure  3b displays 2d plot of $P_{\text{b}}$ versus $f$ at different powers. At resonance $f_{0}$, the heating power $P_{\text{b}}$ grows by increasing input power due to the rising of average photon number in the resonator. The thermal measurement results in a highly frequency independent reference level of the Lorentzian absorption signal $P_{\text{b}}$.
Figure 6: High frequency mode observation, above 20 GHz. a, $P_{\text{b}}$ at around resonance of the mode $3f_{0}~{}\sim 21.2~{}\mathrm{GHz}$. With the bolometer, we can clearly observe the Lorentzian power $P_{\textrm{b}}$ signal around the resonance $3f_{0}$. b, Due to the limitation of the bandwidth of the low-temperature amplifier and isolators, the $S_{21}$ measurement with VNA cannot detect the mode of $3f_{0}$.
@@
We normalize the measured $P_{\text{b}}$ by its magnitude at resonance $P_{\text{b}}(f)/P_{\text{b}}(f_{0})$ as plotted in Fig  4 for $P_{\text{in}}=-110.9~{}\mathrm{dBm}$. We fit the data with normalized Lorentzian function of model Eq.1 to obtain the total loss rate $\gamma_{\text{t}}$ as shown in Fig  4. The plot of $\gamma_{\text{t}}$ versus $P_{\text{in}}$ is shown in Fig  5a. It is known that $\gamma_{\text{t}}$ has power dependence that is due to varying internal loss rate $\gamma_{\text{i}}$. We observe two different behaviors at below and above power around $P_{\text{in}}=-102~{}\mathrm{dBm}$. To understand this behavior, we estimate the internal quality factor of the resonator $Q_{\text{i}}$ as plotted in Fig  5b calculated by Eq. 2. The increase of $Q_{\text{i}}$ at low power regime is well known to be due to saturation of TLSs [19], and at sufficiently high power the resonator is heated up [20]. We confirm this by comparing the data with model of internal quality factor due to TLSs and QPs.
@@
where $Q_{\text{i},\text{TLS}}=\sqrt{1+(P/P_{c})^{\beta/2}}/\delta^{0}_{\text{TLS}}\tanh{(hf_{0}/2k_{B}T})$ is due to losses by TLSs [19]and $Q_{\text{i},\text{QPS}}=Ae^{-P/P_{\text{q}}}$ is phenomenological model describing quasiparticle losses [20]. The list of values of fitting parameters used is shown in Table 1.
Furthermore, we measure the $S_{21}$ and $P_{\text{b}}$ at frequency above 20 GHz to detect the mode of $3f_{0}$ of the resonator shown in Fig  6a-b. With the bolometer, we observe the Lorentzian heating at around frequency $f~{}\sim 21.2~{}\mathrm{GHz}$, which is the mode of $3f_{0}$. While with the $S_{21}$ scattering measurement, due to the limitation of the bandwidth of the rf low-temperature amplifier and rf isolators, we cannot detect any signal at this frequency range. The bolometer has an estimated cut-off frequency of about $200~{}\mathrm{GHz}$, limited by the LR-circuit resonance cutoff $f_{\text{c}}=R_{\text{b}}/2\pi L_{\text{b}}$, where $L_{\text{b}}\sim 8.5~{}\mathrm{pH}$ is the inductance of the Cu wire. This is a clear demonstration of the advantage of the bolometer thanks to its broad operational frequency compared to standard rf measurement scheme. Moreover, there is a regime where the thermal spectrometer is fully calibration-free. When the temperature difference $\Delta T=T_{\text{b}}-T_{0}$ is small ($|\Delta T/T_{\text{0}}|\ll 1$), we can approximate Eq.3 by $P_{\text{b}}\approx 5\Sigma VT^{4}_{0}\Delta T$ [35]. Since in this regime $T_{\text{b}}\sim V_{\text{th}}$ is approximately linear, we have $P_{b}\sim V_{\text{th}}$. Therefore, $V_{\text{th}}$ follows already the Lorentzian spectrum. In this case, we can fit directly $V_{\text{th}}$ with Lorentzian function as shown in Fig  7a. Figure  7b displays $\gamma_{\text{t}}$ obtained from fitting $P_{\text{b}}$ and $V_{\text{th}}$. They show equal results at low power but deviate at higher power levels due to large temperature rise, i.e. non-linearity in Eq.3.

### v2

@@ Introduction
Figure 1: Principle of characterizing a resonator with a bolometer. a, Photons in the resonator (dark blue) are excited due to the input microwave tone $P_{\textrm{in}}$ that is injected to the nearby feedline (light blue). The photons escape via several loss channels, a fraction of them to a nearby capacitively coupled bolometer consisting of an absorber (red) and thermometer. By measuring temperature rise in the bolometer due to this photon leakage, the total loss rate $2\pi\gamma_{\textrm{t}}=2\pi(\gamma_{\textrm{f}}+\gamma_{\textrm{i}}+\gamma_{\textrm{b}})$ and resonance frequency of the resonator ($f_{0}$) can be determined. b, The system is the case of a driven resonator (a harmonic oscillator) coupled to several thermal baths that absorb the photons from the resonator. c, Equivalent lumped circuit. The bolometer is heated up due to the power $P_{\textrm{b}}$ transmitted to the absorber $R_{\textrm{b}}$. The magnitude of $P_{\textrm{b}}$ depends on the impedance of the other circuit elements $R,L,{\textrm{and}}~{}C$. $P_{\textrm{i}}$ is the power dissipated in the internal bath.
Detection in superconducting devices is a central key in the characterization and operation of the devices. For example, RF spectroscopy becomes a standard technique to measure qubit energy spectrum and its coherent interaction with microwave photons [16, 17, 18]. The study of a resonator and its loss mechanisms relies on the spectroscopy technique, either transmission ($S_{21}$) or reflection ($S_{11}$) measurement [19, 20]. Conventional RF spectroscopy, involving amplification of a microwave probe tone at different temperature stages followed by heterodyne detection at room temperature, requires electronics and cryogenic elements such as vector network analyzer (VNA), low-temperature amplifiers, circulators, and superconducting coaxial lines. Precise cabling, grounding, and connection of the lines and components are crucial to avoid any microwave mismatch and leakage, that can result to parasitic resonances and suppression of the desired signals [21].
@@
A bolometer consists of an energy absorbing element and the associated thermometer detecting temperature changes due to energy input [22, 23]. Bolometric detection in a superconducting circuit offers advantages compared to conventional schemes of RF measurement. Using a bolometric detector, that can be placed at the millikelvin stage, the use of cryogenic elements can be avoided. A single-shot qubit readout has been demonstrated by using a bolometer that is located on a separate chip [24]. Without the use of low-temperature parametric amplifier, the authors could measure the qubit state by monitoring temperature in the bolometer. Another advantage is the broadband operation of the bolometer, while a typical RF setup is limited to narrow frequency range of $4-8~{}\textrm{GHz}$. A nano-bolometer has been used to calibrate the attenuation of a coaxial line in a broad frequency range [25]. A bolometric measurement of a high frequency Josephson ac current has been demonstrated [26] where the authors could detect ac current up to 100 GHz. Moreover, the bolometer could reach a high energy resolution and therefore can measure small energies as demonstrated by proximitized thermometer and graphene as the thermal detector [27, 28], and it is a potential candidate of a single photon detector.
@@
The principle of characterizing a superconducting resonator by an on-chip bolometer is shown in Fig. 1a. The resonator (dark blue) is coupled to a feedline (light blue) via capacitance $C_{\textrm{f}}$. The open-circuit end of the resonator (left) is coupled by capacitance $C_{\textrm{b}}$ to a bolometer that consists of an absorber (red) with resistance $R_{\textrm{b}}$ and the associated thermometer to detect temperature change due to input energy. A microwave tone $P_{\textrm{in}}$ is injected through the input feedline to excite photons in the resonator. The photons subsequently decay via several loss channels. Internal loss is mostly due to TLSs or QPs depending on power and temperature, and external losses are dominated by the proximity of the feedline and the bolometer itself. Thus, the total loss rate is $2\pi\gamma_{\textrm{t}}=2\pi(\gamma_{\textrm{b}}+\gamma_{\textrm{i}}+\gamma_{\textrm{f}})$, where $2\pi\gamma_{\textrm{i}}$ is the internal loss rate, $2\pi\gamma_{\textrm{f}}$ is loss rate to the feedline, and $2\pi\gamma_{\textrm{b}}$ is loss rate to the bolometer. The system is the case of a driven resonator (a harmonic oscillator) coupled to several thermal baths, as shown in Fig.1b. The photons that are absorbed by the bolometer heat the resistor with power $P_{\textrm{b}}$, leading to temperature rise $T_{\textrm{b}}$. The power $P_{\textrm{b}}$ takes Lorentzian form as a function of the driving frequency around resonance $f_{0}$ with the linewidth $\gamma_{\textrm{t}}$. Due to the linearity of the resonator, $P_{\textrm{b}}$ is independent on bolometer temperature and takes a form
$P_{\mathrm{b}}(f)=\frac{1}{2}\frac{\gamma_{\textrm{f}}\gamma_{\textrm{b}}}{(f-f_{0})^{2}+(\gamma_{\textrm{t}}/2)^{2}}P_{\mathrm{in}},$ (1)
a result that one can obtain both by circuit theory and open quantum system approach (see Supplementary Information I and II). By measuring the frequency dependence of $P_{\textrm{b}}$, which can be measured by a steady-state temperature measurement of the bolometer, the total-loss rate $2\pi\gamma_{\textrm{t}}$ and the resonance frequency $f_{0}$ of the resonator can thus be determined.
@@
Figure 3: Power measured by bolometer. a, Heating power $P_{\text{b}}$ and bolometer temperature $T_{\text{b}}$ due to photon leakage from the resonator. b, $P_{\text{b}}$ versus $f$ at different input powers varying from $-130~{}\mathrm{dBm}$ to $-95~{}\mathrm{dBm}$. Top axis shows the average photon number $N$ in the resonator.
@@
Figure 5: Linewidth and quality factor. a, Total linewidth $\gamma_{\textrm{t}}$ and total quality factor $Q_{\textrm{t}}$ versus $P_{\text{in}}$. The error bars are uncertainty from the fitting. b, Comparison between measured internal quality factor $Q_{\textrm{i}}$ (orange) and TLSs $Q_{\textrm{i,TLS}}$ (blue) and QPS $Q_{\textrm{i,QPS}}$ (red) model. The grey area shows total uncertainty due to additional potential error in simulated circuit parameters.
@@
The studied device consists of a quarter-wavelength ($\lambda/4$) coplanar waveguide (CPW) resonator capacitively probed by the feedline and the on-chip bolometer as shown in the device layout in Fig. 2(a). Both resonator and feedline are made of niobium (Nb) film on top of silicon (Si) substrate with AlOx sublayer. We design $C_{\mathrm{f}}\sim 13.85~{}\mathrm{fF}$, $C_{\mathrm{b}}\sim 19.6~{}\mathrm{fF}$, $R_{\mathrm{b}}$ is about $12.23~{}\Omega$, equivalent inductance of the resonator is $L=1.44~{}\mathrm{nH}$, and equivalent capacitance of the resonator is $C=356~{}\mathrm{fF}$. These values correspond to the quality factor of the feedline and bolometer as $Q_{\mathrm{f}}\sim 1681$ and $Q_{\mathrm{b}}\sim 1716$ respectively. The shorted end of the resonator (right) is shunted by inactive flux qubit consisting of a parallel aluminium (Al) line and Josephson junctions, tuned at zero magnetic flux thus effectively shunted only by the Al line [30, 31]. The open-circuit end of the resonator is capacitively connected to a copper (Cu) film functioning as the bolometer as shown in Fig. 2(b). In the experiment, we employ a pair of normal metal-insulator-superconductor (NIS) junctions in SINIS configuration to probe the temperature ($T_{\textrm{b}}$) in the Cu absorber (red), as shown in Fig. 2(c). The temperature rise manifests in the change of voltage $V_{\textrm{th}}$ accross the SINIS pair [23, 32]. Another single NIS junction is connected to a voltage source ($V_{\textrm{h}}$) for heating the resistor (increasing $T_{\textrm{b}}$) to test and calibrate the thermometry. Once the temperature is determined from the calibration conversion (see Supplementary Information III), the power can be calculated by electron phonon relation as
@@
Figure 6: High frequency mode observation, above 20 GHz. $P_{\text{b}}$ at around resonance of the mode $3f_{0}~{}\sim 21.2~{}\mathrm{GHz}$. With the bolometer, we can clearly observe the Lorentzian power $P_{\textrm{b}}$ signal around the resonance $3f_{0}$ with a linewidth $\sim 80~{}\text{MHz}$. Due to the limitation of the bandwidth of the low-temperature amplifier and isolators, the $S_{21}$ measurement with VNA cannot detect the mode of $3f_{0}$.
@@
We convert the measured $V_{\text{th}}$ to its corresponding temperature $T_{\text{b}}$ and power dissipation $P_{\text{b}}$. By using the conversion formula obtained from the bolometer calibration (see Supplementary Information III), we obtain $T_{\text{b}}$ around the resonance as shown in Fig  3a on the right axis. Although the measurement is carried out at 50 mK, $T_{\text{b}}$ saturates at around 130 mK. This saturation is dominated by heating from the environment with effective background power $P_{\text{e}}\sim 2~{}\text{fW}$ which mostly comes from the DC lines that connect to the NIS junctions [32, 34]. The heating power $P_{\text{b}}$, converted from Eq.3, is plotted in Fig  3a on the left axis. Figure  3b displays 2d plot of $P_{\text{b}}$ versus $f$ at different powers. At resonance $f_{0}$, the heating power $P_{\text{b}}$ grows by increasing input power due to the rising of average photon number in the resonator, $N=P_{\text{b}}(f_{0})Q_{\text{b}}/2\pi hf_{0}^{2}$ (see Supplementary Information IV). The thermal measurement results in a highly frequency independent reference level of the Lorentzian absorption signal $P_{\text{b}}$.
@@
We normalize the measured $P_{\text{b}}$ by its magnitude at resonance $P_{\text{b}}(f)/P_{\text{b}}(f_{0})$ as plotted in Fig  4 for $P_{\text{in}}=-110.9~{}\mathrm{dBm}$. We fit the data with normalized Lorentzian function of model Eq.1 to obtain the total linewidth $\gamma_{\text{t}}$ as shown in Fig  4. The plot of $\gamma_{\text{t}}$ and $Q_{\text{t}}$ versus $P_{\text{in}}$ is shown in Fig 5a. It is known that $\gamma_{\text{t}}$ has power dependence that is due to varying internal loss rate $2\pi\gamma_{\text{i}}$. We observe two different behaviors at below and above power around $P_{\text{in}}=-102~{}\mathrm{dBm}$. To understand this behavior, we estimate the internal quality factor of the resonator $Q_{\text{i}}$ as plotted in Fig  5b calculated by Eq. 2. Extraction of $Q_{\text{i}}$ relies on knowledge of circuit parameters. From measured $f_{\text{0}}$, the equivalent capacitance and inductance are determined by $C=1/8Z_{0}f_{\text{0}}$ and $L=2Z_{0}/\pi^{2}f_{\text{0}}$ (see Supplementary Information I), and the resistance $R_{\text{b}}$ is measured. The $C_{\text{b}}$ and $C_{\text{f}}$ are obtained from simulations. In Fig  5b, the error bars come from the fitting uncertainty, while the gray area covers total uncertainty due to additional potential 2% error in $C_{\text{b}}$ and $C_{\text{f}}$. The increase of $Q_{\text{i}}$ at low power regime is well known to be due to saturation of TLSs [19], and at sufficiently high power the resonator is heated up [20]. We confirm this by comparing the data with model of internal quality factor due to TLSs and QPs.
@@
where $Q_{\text{i},\text{TLS}}=\sqrt{1+(P/P_{c})^{\beta/2}}/\delta^{0}_{\text{TLS}}\tanh{(hf_{0}/2k_{B}T_{0}})$ is due to losses by TLSs [19]and $Q_{\text{i},\text{QPS}}=Ae^{-P/P_{\text{q}}}$ is phenomenological model describing quasiparticle losses [20]. The list of values of fitting parameters used is shown in Table 1.
Furthermore, we measure the $S_{21}$ and $P_{\text{b}}$ at frequency above 20 GHz to detect the mode of $3f_{0}$ of the resonator. With the bolometer, we observe the Lorentzian heating at around frequency $f~{}\sim 21.2~{}\mathrm{GHz}$, which is the mode of $3f_{0}$ as shown in Fig. 6. The extracted linewidth is around $\sim 80~{}\text{MHz}$. While with the $S_{21}$ scattering measurement, due to the limitation of the bandwidth of the RF low-temperature amplifier and RF isolators, we cannot detect any signal at this frequency range. The bolometer has an estimated cut-off frequency of about $200~{}\mathrm{GHz}$, limited by the LR-circuit resonance cutoff $f_{\text{c}}=R_{\text{b}}/2\pi L_{\text{b}}$, where $L_{\text{b}}\sim 8.5~{}\mathrm{pH}$ is the inductance of the Cu wire. This is a clear demonstration of the advantage of the bolometer thanks to its broad operational frequency compared to standard RF measurement scheme.
Moreover, there is a regime where the thermal spectrometer is fully calibration-free. When the temperature difference $\Delta T=T_{\text{b}}-T_{0}$ is small ($|\Delta T/T_{\text{0}}|\ll 1$), we can approximate Eq.3 by $P_{\text{b}}\approx 5\Sigma\Omega T^{4}_{0}\Delta T$ [35]. Since in this regime $T_{\text{b}}\sim V_{\text{th}}$ is approximately linear, we have $P_{b}\sim V_{\text{th}}$. Therefore, $V_{\text{th}}$ follows already the Lorentzian spectrum. In this case, we can fit directly $V_{\text{th}}$ with Lorentzian function without the need to convert it to $P_{\text{b}}$ as shown in Fig  7a. Figure  7b displays $\gamma_{\text{t}}$ obtained from fitting $P_{\text{b}}$ and $V_{\text{th}}$. They show equal results at low power, nearby single-photon regime, but deviate at higher power levels due to large temperature rise, i.e. non-linearity in Eq.3.